# Stitched ViTs are Flexible Vision Backbones

### v1

@@ 5 Experiments
In this section, we first show the advantage of SN-Netv2 over SN-Netv1 on ImageNet-1K [38]. Next, we conduct experiments on downstream dense prediction tasks to show the strong performance of SN-Netv2 as a promising flexible vision backbone, including semantic segmentation on ADE20K [57], COCO-Stuff [4], depth estimation on NYUv2 [40] and object detection on COCO-2017 [29].
@@ 5.1 ImageNet Classification
Implementation details. We experiment with three combinations of stitching for DeiT3 [45], namely DeiT3-S/B, DeiT3-S/L and DeiT3-B/L. We train each setting on ImageNet-1K by 30 epochs with a total batch size of 256 on 8 V100 GPUs. The learning rate is 0.1$\times$ to that of default training time learning rate in DeiT3. During training, anchors adopt the same stochastic layer drop rate as the pretrained model family, i.e., 0.05, 0.15, 0.4 for DeiT3-S, DeiT3-B, DeiT3-L, respectively. All other hyperparameters are the same as in DeiT3. For all experiments, we adopt a rank of 16 for LoRA SL and 100 images for its LS initialization. We evaluate the performance by Top-1 accuracy (%).
Figure 6: Semantic segmentation results of SN-Netv2 on COCO-Stuff-10K by stitching DeiT3-S/L and DeiT3-B/L. We highlight the stitches at the Pareto frontier in blue line.
@@ 5.2 Semantic Segmentation
ADE20K results. We report our ADE20K results in Figure 5. Specifically, based on DeiT3-S and DeiT3-L, SN-Netv2 demonstrates strong performance against anchor performance while simultaneously supporting a diverse range of resource constraints. By stitching DeiT3-B and DeiT3-L, SN-Netv2 achieves equal performance with anchors at their FLOPs. Moreover, by ploting the stitches on Pareto Frontier, we show SN-Netv2 smoothly interpolates the performance of two solo backbone settings, which strongly demonstrates the effectiveness of our method.
COCO-Stuff-10K results. In Figure 6, we show that stitching DeiT3-S and DeiT3-L under SN-Netv2 even achieves better performance than anchors. More impressively, by stitching DeiT3-B and DeiT3-L, we found some stitches that achieve better performance than the large anchor at a lower FLOPs. It implies that the original plain ViTs may not be the best architecture in different domains.
@@
Figure 7: Results of stitching DeiT3-S and DeiT3-L on NYUv2 under the framework of DPT.
Figure 8: Results of stitching DeiT3-B and DeiT3-L on NYUv2 under the framework of DPT.
@@ 5.3 Depth Estimation
Implementation details. Based on DPT [37], we conduct experiments on NYUv2 [40] dataset and train SN-Netv2 by stitching DeiT3-S/L and DeiT3-B/L. Specificially, we train each model on 4 A100 GPUs with a total batch size of 16. We set the training epochs to 24. We adopt AdamW optimizer with an initial learning rate of $2\times 10^{-5}$. For LoRA SL, we use a rank of 4 for all experiments. We utilize common metrics for evaluting the performance on depth estimation, including $\delta>1.25$, $\delta>1.25^{2}$, $\delta>1.25^{3}$ (where higher values indicate better performance), as well as AbsRel, RMSE, and Log10 (where lower values indicate better performance).
@@
Results. We report our experiment results on NYUv2 [40] in Figure 7 and Figure 8. Overall, SN-Netv2 demonstrates highly competitive performance compared to the anchor models, while simultaneously achieving a smooth performance-efficiency curve. In particular, we found stitching DeiT3-B and DeiT3-L is slightly better, which we assume a smaller gap in model complexity between the anchors can help to achieve better performance, which aligns with the observations of SN-Netv1 [33].
5.4 Object Detection and Instance Segmentation
Implementation details. We experiment SN-Netv2 on COCO-2017 and adopt Mask R-CNN based ViTDet [28]. We train all models including individual anchors on 8 A100 GPUs with a total batch size of 8 for 100 epochs. We set the rank as 16 for LoRA SL. Besides, we adopt the same layer decay rate as the baselines for different anchors. All other hyperparameters adopt the default setting in detectron2 [47].
Results. Based on DeiT3-S and DeiT3-L, we report the performance of SN-Netv2 on object detection and instance segmentation in Figure 9. Overall, we found SN-Netv2 exhibits strong flexibility on the detection task as well, as evidenced by the smooth metrics under various resource constraints. This again demonstrates that SN-Netv2 can serve as a flexible backbone on a wide range of CV tasks. In particular, under the similar training cost ($\sim$1500 GPU hours), SN-Netv2 can achieve comparable performance at the FLOPs of individually trained DeiT3-L (51.2 vs. 51.9 bbox AP), while supporting many FLOPs-accuracy trade-offs at runtime without additional training cost. However, we observe a performance gap between individually trained DeiT3-S with that of SN-Netv2 at the same FLOPs (42.9 vs. 46.6 bbox AP). For this, we hypothesize that with the heavy decoder in ViTDet (27×27$\times$27 × larger than SETR decoder), simultaneously ensuring the performance of hundreds of stitches as backbones can be more difficult. Nevertheless, SN-Netv2 still achieves superior advantage in terms of the training efficiency as it only requires training once while covering many FLOPs-accuracy trade-offs. We leave the improvement for future work.
@@ 5.4 Object Detection and Instance Segmentation
5.5 Ablation Study
Effect of improved techniques. Based on DeiT3-S and DeiT3-L, we conduct experiments on ImageNet-1K to show the effect of Two-way stitching, resource-constrained sampling and LoRA stitching layers. As shown in Figure 10, benefiting form Two-way stitching, SN-Net successfully finds better stitching configurations at the relatively low FLOPs constraints except for the small anchor where it drops significantly. However, with the help of ROS, we improve the performance of DeiT3-S and therefore ensure the overall performance curve. Finally, we show LoRA SL can achieve comparable performance with the fully finetuned baseline in Figure 10 (c).
Effect of LoRA SL on downstream tasks. We have shown that applying SN-Netv2 with or without LoRA SL on ImageNet-1K has a minor effect on the performance. However, we emphasize that low-rank adaptation for stitching layers is critical for ensuring good performance on downstream tasks. In Figure 11, we report the results of SN-Netv2 with/without applying LoRA to stitching layers on ADE20K. As it shows, without LoRA SL, there is a noticeable performance drop for anchors when stitching DeiT3-S/L. The issue is even more pronounced when stitching DeiT3-B/L, resulting in a highly unstable performance curve. However, after applying LoRA to stitching layer, we achieve a more stable and better performance curve. In this case, we speculate that low-rank update of the stitching layer can stabilize anchor learning, thus ensuring good performance of intermediate stitches in SN-Netv2. We show the effect of different ranks in Section A.4.
Figure 11: Performance comparison of SN-Netv2 with (Figure (b) and (d)) and without (Figure (a) and (c)) low-rank adaptation for stitching layers on ADE20K.
Figure 12: Figure (a): Comparing parameter-efficient tuning (PET) with fully finetuning (Full) on SN-Netv2. Figure (b): Benchmarking different pretrained weights of plain ViTs under SN-Netv2. Experiments are conducted on COCO-Stuff-10K and based on Base/Large variants of ViTs.
Compared to parameter-efficient tuning on backbone. We investigate the impact of parameter-efficient tuning (PET) on the performance of SN-Netv2 by stithcing DeiT3-B/L. Specifically, we apply LoRA with a rank of 4 to the backbone and train it using the same schedule as fully-finetuning on COCO-Stuff-10K. As shown in Figure 12 (a), SN-Netv2 can produce good performance-efficiency trade-offs under PET as well. Moreover, it even achieves better performance than individually trained anchors with PET (e.g., 44.8 vs. 45.8 with DeiT3-L), for which we conjecture SN-Netv2 may regularize the large anchor learning when combining with PET.
Benchmarking other pretrained ViT weights. In Figure 12 (b), we report the results of stitching different pretrained weights based on the base and large variants of ViTs, including MAE [20], SAM [26], AugReg [41] and BEiTv2 [34]. Overall, SN-Netv2 consistently generates good Pareto frontiers for these weights. As pre-training objectives and datasets differ across ViTs, our results in Figure 12 (b) demonstrate varying performance when stitching different ViT weights, where DeiT3 achieves the best performance. Therefore, we choose DeiT3 as the default weights.
6 Conclusion and Future Work
We have introduced SN-Netv2, a systematically improved training framework that effectively employs off-the-shelf foundation ViTs to obtain a flexible and strong vision backbone on downstream dense prediction tasks. In particular, we have first proposed a Two-way stitching strategy to enlarge the stitching space and effectively identified more optimal stitches in certain resource constraints. Next, we have devised a resource-constrained sampling approach to ensure balanced training for stitches that reside on different resource constraints, thereby enhancing the overall performance of SN-Netv2. Furthermore, we have demonstrated that low-rank adaptation of stitching layers yields competitive performance on ImageNet-1K while playing a crucial role in stabilizing training on downstream tasks. Extensive experiments across various datasets have demonstrated that SN-Netv2 achieves great training efficiency and inference flexibility than typical ViT backbone adoption, benefiting the massive model deployment in real-world applications.
Limitations and societal impact. This paper mainly aims to improve SN-Netv1 for flexible ViT adoption on downstream tasks, which however leaves parameter efficient approaches [24, 23, 7] under-explored. Future work may also consider a better training strategy to improve the performance of stitches on the Pareto frontier. Besides, SN-Netv2 requires multi-GPU training, which inevitably results in a substantial electricity consumption and carbon emission.
Appendix A Appendix
We organize our supplementary material as follows.
@@ Appendix A Appendix
  • In Section A.1, we compare different stitching types at initialization based on ImageNet-1K.
@@
  • In Section A.2, we compare different stitching types after training based on ImageNet-1K.
  • In Section A.3, we compare SN-Netv2 with SN-Netv1 on semantic segmentation.
  • In Section A.4, we explore the effect of different ranks in LoRA SL.
  • In Section A.5, we describe the details of the benchmarked pretrained ViTs.
@@
A.1 Performance Comparisons of Different Stitching Types at Initialization
@@
By default, we randomly sample 100 images from the train set and solve the least-square problem to initialize the stitching layers. After initialization, we directly compare the performance of different types of stitches based on DeiT3 models and ImageNet-1K. Overall, the observations depicted in Figure 13 indicate that Fast-to-Slow (FS) outperforms at high FLOPs, while Slow-to-Fast (SF) excels at low FLOPs. This highlights the effectiveness of Two-way stitching, as no single stitching direction emerges as the superior choice across all FLOPs intervals. Similarly, in Figure 14, we find that Fast-Slow-Fast (FSF) generally outperforms Slow-Fast-Slow (SFS) at low FLOPs, whereas SFS stitching prevails at high FLOPs. This implies that benefiting from the enlarged stitching space, SN-Netv2 can find a large number of more optimal architectures than SN-Netv1 (which uses FS only), enabling them to achieve favorable performance right from the initialization phase.
Figure 16: Performance comparison between SN-Netv1 and SN-Netv2 on ADE20K and COCO-Stuff-10K by stitching DeiT3-S and DeiT3-L.
@@
A.2 Performance Comparison of Different Stitching Types after Training
@@
We show in Figure 15 that different stitching types present varying performance in different resource constraints after training. Similar to the phenomenon in Section A.1, our findings reveal that SF outperforms FS at low FLOPs, while FS becomes superior at high FLOPs, particularly for ImageNet-1K and ADE20K. We attribute this to the fact that networks adopting a few early blocks from the other anchor tend to perform better, indicating that stitching at the early stages of ViTs is more effective when stitched only once. For this, we assume early blocks involve with general low level features [22, 48], which makes them more amenable to stitch. Besides, FSF stitches are generally better than that of SFS after training. These observations can serve as valuable guidelines for future stitching designs.
A.3 Performance Comparison of SN-Netv1 and SN-Netv2 on Semantic Segmentation
Based on DeiT3-S and DeiT3-L, we compare the performance of SN-Netv2 to SN-Netv1 on ADE20K and COCO-Stuff-10K. By default, we adopt a rank of 16 for LoRA SL and train each model on ADE20K with a 160K schedule and COCO-Stuff-10K with a 80K schedule. The total batch size is 16 for all experiments. As shown in Figure 16, it is evident that SN-Netv2 exhibits significant improvement for stitches at low FLOPs, while also achieving competitive performance at high FLOPs. Notably, the highlighted Pareto frontier demonstrates that SN-Netv2 surpasses SN-Netv1 in terms of the overall performance-efficiency trade-offs.
A.4 Effect of Different Ranks in LoRA SL
In Figure 17, we explore the effect of different ranks in LoRA SL by stitching DeiT3-S and DeiT3-L on COCO-Stuff-10K. In general, we observe that different low ranks perform similar, where they can effectively produce smoothly increasing Pareto frontiers. However, without the low-rank update, the performance of the stitches at the higher FLOPs drops (>250G FLOPs). Therefore, it indicates that low-rank update may regularize stitches learning during the training time.
A.5 Details of Benchmarked Pretrained ViTs
We have shown in Section 5.5 that different pretrained ViTs under SN-Netv2 performs differently due to the various training objectives and datasets. In this section, we provide the details of the benchmarked ViT weights.

### v2

@@ 5 Experiments
In this section, we first show the advantage of SN-Netv2 over SN-Netv1 on ImageNet-1K [38]. Next, we conduct experiments on downstream dense prediction tasks to show the strong performance of SN-Netv2 as a promising flexible vision backbone, including semantic segmentation on ADE20K [57], COCO-Stuff [4], and depth estimation on NYUv2 [40]. Due to the limited space, we provide more experiments on COCO detection in the supplement.
@@ 5.1 ImageNet Classification
Implementation details. We experiment with three combinations of stitching for DeiT3 [45], namely DeiT3-S/B, DeiT3-S/L and DeiT3-B/L. We train each setting on ImageNet-1K by 50 epochs with a total batch size of 256 on 8 V100 GPUs. The learning rate is 0.1$\times$ to that of default training time learning rate in DeiT3. During training, anchors adopt the same stochastic layer drop rate as the pretrained model family, i.e., 0.05, 0.15, 0.4 for DeiT3-S, DeiT3-B, DeiT3-L, respectively. All other hyperparameters are the same as in DeiT3. For all experiments, we adopt a rank of 16 for LoRA SL and 100 images for its Least-square initialization [33]. We evaluate the performance by Top-1 accuracy (%).
@@ 5.2 Semantic Segmentation
Figure 10: Performance comparison of SN-Netv2 with (Figures (b) and (d)) and without (Figures (a) and (c)) low-rank adaptation for stitching layers on ADE20K.
ADE20K results. We report our ADE20K results in Figure 8. Specifically, based on DeiT3-S and DeiT3-L, SN-Netv2 demonstrates strong performance against anchor performance while simultaneously supporting a diverse range of resource constraints. By stitching DeiT3-B and DeiT3-L, SN-Netv2 achieves equal performance with anchors at their FLOPs. Moreover, by ploting the stitches on Pareto Frontier, we show SN-Netv2 smoothly interpolates the performance of two solo backbone settings. Compared to SN-Netv1, SN-Netv2 obtains a better overall Pareto frontier, demonstrating the advantage of our method.
COCO-Stuff-10K results. In Figure 8, we show that stitching DeiT3-S and DeiT3-L under SN-Netv2 even achieves better performance than anchors. More impressively, by stitching DeiT3-B and DeiT3-L, we found some stitches that achieve better performance than the large anchor at a lower FLOPs. It implies that the original plain ViTs may not be the best architecture in different domains. Finally, we can observe consistently better frontiers when comparing with SN-Netv1.
@@ 5.3 Depth Estimation
Based on DPT [37], we conduct experiments on NYUv2 [40] dataset and train SN-Netv2 by stitching DeiT3-S/L. We report the experiment results in Figure 9. Overall, SN-Netv2 demonstrates highly competitive performance compared to the anchor models, while simultaneously achieving a better overall performance-efficiency curve than SN-Netv1 across different metrics. We provide implementation details in the supplementary material.
Table 2: Ablation study of each component in SN-Netv2 based on ADE20K by stitching DeiT3-S and DeiT3-L. Results are shown in mIoU. The FLOPs constraints at 34G and 363G indicate the same architecture as the DeiT3-S and DeiT3-L, respectively.
FLOPs Constraints 34G 61G 88G 116G 143G 171G 198G 225G 253G 280G 308G 335G 363G
SN-Netv1 40.4 41.5 43.2 44.2 45.2 46.4 47.9 49.2 50.0 51.0 51.4 51.2 51.6
SN-Netv1 + TWS 42.8 43.3 44.2 45.4 46.5 47.8 48.5 49.3 49.6 49.6 49.8 50.2 50.5
SN-Netv1 + TWS + ROS 43.7 43.9 45.5 46.7 47.5 48.1 48.8 49.3 49.8 50.0 50.2 50.3 50.7
SN-Netv1 + TWS + ROS + LoRA SL 45.6 45.7 46.8 47.7 48.1 49.0 48.9 49.4 49.3 49.9 50.7 50.8 51.4
Table 3: Comparison between full and LoRA finetunings for the entire backbone on COCO-Stuff10K. We measure the training speed in “s/iter”, which is the time cost in seconds for each training iteration on 4 A100. FLOPs and mIoU in SN-Netv2 are represented as a range.
Method FLOPs (G) mIoU s/iter
DeiT3-L 363 47.8 0.95
SN-Netv2 34 - 363 40.1 - 48.2 0.68
DeiT3-L + LoRA 363 44.8 0.90
SN-Netv2 + LoRA 34 - 363 42.6 - 45.8 0.65
5.4 Ablation Study
Effect of the three new components. Based on DeiT3-S and DeiT3-L, we conduct experiments on ADE20K and show the ablation results in Table 2. First, we can see that, benefiting from TWS, SN-Netv2 finds better stitching configurations at the relatively low FLOPs (33G-198G). However, there is a performance gap at the FLOPs of DeiT3-S when compared with the individually trained one (45.5 vs. 42.8 mIoU). On ImageNet, the performance drop is more significant (-17.2% Top-1 accuracy, shown in the supplementary material). Second, the proposed ROS reduces the performance gap at the FLOPs of the small anchor and improves the overall performance curve. Finally, LoRA SL helps to further improve the performance significantly, achieving comparable performance with the individually trained DeiT3-S (45.6 vs 45.5 mIoU), while maintaining competitive performance at high FLOPs.
Effect of LoRA SL for stabilizing training. In our experiments, we empirically found that directly enlarging the space with TWS and ROS on downstream tasks results in a sub-optimal Pareto frontier, see Figure 10. To this end, the proposed low-rank adaptation for stitching layers is critical for ensuring the overall good performance on downstream tasks. As shown in Figure 10, without LoRA SL, there is a noticeable performance drop at anchors when stitching DeiT3-S/L. The issue is even more pronounced when stitching DeiT3-B/L, resulting in a highly unstable performance curve. On the contrary, after applying LoRA to stitching layers, we achieve a more stable and better performance curve. We speculate that low-rank updates of stitching layers can stabilize anchor learning, thus ensuring good performance of intermediate stitches in SN-Netv2. We show the effect of different ranks in the supplementary material.
Compare with parameter-efficient finetuning on the backbone. SN-Netv2 is compatible with existing parameter-efficient finetuning (PEFT) techniques (e.g., LoRA [23]) on the entire backbone to achieve faster training speed and parameter efficiency. Here, we conduct experiments on COCO-Stuff-10K, and report the results in Table 3. First, we can see that SN-Netv2 can achieve faster training speed than the individual large ViT training while obtaining a wide range of trade-offs. This is because during training SN-Netv2 selects a stitch from different FLOPs constraints, which is more efficient than always training the same largest ViT. Second, compared to the individual large ViT training with LoRA, SN-Netv2 with LoRA achieves even better performance (45.8 vs. 44.8), for which we conjecture SN-Netv2 may regularize the large anchor learning when combining with PEFT [17], besides faster training speed and more trade-offs.
6 Conclusion and Future Work
We have introduced SN-Netv2, a systematically improved training framework that effectively employs off-the-shelf foundation ViTs to obtain a flexible and strong vision backbone for downstream dense prediction tasks. Particularly, we proposed a two-way stitching strategy to enlarge the stitching space, devised a resource-constrained sampling approach to ensure balanced training for stitches that reside on different resource constraints, and introduced low-rank adaptation of stitching layers for stabilizing training on downstream tasks. Extensive experiments across various datasets have demonstrated that SN-Netv2 achieves great training efficiency and inference flexibility than typical ViT backbone adoption, benefiting flexible model deployments in diverse real-world applications.
Limitations. This work mainly aims to improve SN-Netv1 for flexible ViT adoption on downstream dense prediction tasks, which however leaves parameter efficient approaches [24, 23, 7] under-explored. Future work may also consider a better training strategy to improve the performance of stitches on the Pareto frontier.
Appendix
@@ 6 Conclusion and Future Work
We organize our supplementary material as following,
  • In Section A, we ablate our proposed three new components on ImageNet-1K, namely Two-way stitching (TWS), Resource-constrained Sampling (ROS) and LoRA Stitching Layers (LoRA SL).
  • In Section B, we explore the effect of different ranks in LoRA SL.
  • In Section C, we report the implementation details of depth estimation experiments on NYUv2.
@@
  • In Section D, we compare different stitching types at initialization based on ImageNet-1K.
@@
  • In Section E, we compare different stitching types after training based on ImageNet-1K.
  • In Section F, we show the detailed performance curve of applying parameter-efficient finetuning (PEFT) on the entire backbone.
  • In Section G, we study the effect of using different pretrained ViTs in SN-Netv2.
Appendix A Effect of The Three New Components on ImageNet-1K
Based on DeiT3-S and DeiT3-L, we conduct experiments to show the effect of Two-way stitching, resource-constrained sampling and LoRA stitching layers on the pretraining dataset ImageNet-1K. As shown in Figure 11 (b), benefiting form Two-way stitching, SN-Net successfully finds better stitching configurations at the relatively low FLOPs constraints except for the small anchor where it drops significantly. However, with the help of ROS, we improve the performance of DeiT3-S and therefore ensure the overall performance curve in Figure 11 (c). Finally, we show LoRA SL can achieve comparable performance with the fully finetuned baseline on ImageNet-1K in Figure 11 (d).
Appendix B Effect of Different Ranks in LoRA SL
In Figure 12, we explore the effect of different ranks in LoRA SL by stitching DeiT3-S and DeiT3-L on COCO-Stuff-10K. In general, we observe that different low ranks perform similar, where they can effectively produce smoothly increasing Pareto frontiers. However, without the low-rank update, the performance of the stitches at the higher FLOPs drops ($>$250G FLOPs). Therefore, it indicates that the low-rank update may regularize stitches learning during the training time.
@@ Appendix B Effect of Different Ranks in LoRA SL
Appendix C Implementation Details of Depth Estimation Experiments
Based on DPT [37], we conduct experiments on NYUv2 [40] dataset and train SN-Netv2 by stitching DeiT3-S/L and DeiT3-B/L. Specificially, we train each model on 4 A100 GPUs with a total batch size of 16. We set the training epochs to 24. We adopt AdamW optimizer with an initial learning rate of $2\times 10^{-5}$. For LoRA SL, we use a rank of 4 for all experiments. We utilize common metrics for evaluting the performance on depth estimation, including $\delta>1.25$, $\delta>1.25^{2}$, $\delta>1.25^{3}$ (where higher values indicate better performance), as well as AbsRel, RMSE, and Log10 (where lower values indicate better performance). All other hyperparameters are set as the default setting in DPT.
Appendix D Performance Comparisons of Different Stitching Types at Initialization
@@
By default, we randomly sample 100 images from the training set and solve the least-square problem to initialize the stitching layers. After initialization, we directly compare the performance of different types of stitches based on DeiT3 models and ImageNet-1K. Overall, the observations depicted in Figure 13 indicate that Fast-to-Slow (FS) outperforms at high FLOPs, while Slow-to-Fast (SF) excels at low FLOPs. This highlights the effectiveness of Two-way stitching, as no single stitching direction emerges as the superior choice across all FLOPs intervals. Similarly, in Figure 14, we find that Fast-Slow-Fast (FSF) generally outperforms Slow-Fast-Slow (SFS) at low FLOPs, whereas SFS stitching prevails at high FLOPs. This implies that benefiting from the enlarged stitching space, SN-Netv2 can find a large number of more optimal architectures than SN-Netv1 (which uses FS only), enabling them to achieve favorable performance right from the initialization phase.
@@
Appendix E Performance Comparison of Different Stitching Types after Training
@@
We show in Figure 15 that different stitching types present varying performance under different resource constraints after training. Similar to the phenomenon in Section D, our findings reveal that SF outperforms FS at low FLOPs, while FS becomes superior at high FLOPs, particularly for ImageNet-1K and ADE20K. We attribute this to the fact that networks adopting a few early blocks from the other anchors tend to perform better, indicating that stitching at the early stages of ViTs is more effective when stitched only once. For this, we assume early blocks involve with general low level features [22, 48], which makes them more amenable to stitch. Besides, FSF stitches are generally better than that of SFS after training. These observations can serve as valuable guidelines for future stitching designs.
Figure 16: Figure (a): Comparing parameter-efficient tuning (PEFT) with fully finetuning (Full) on SN-Netv2. Figure (b): Benchmarking different pretrained weights of plain ViTs under SN-Netv2. Experiments are conducted on COCO-Stuff-10K and based on Base/Large variants of ViTs.
Appendix F Detailed Performance Comparison for SN-Netv2 with PEFT
We have compared applying LoRA PEFT on the entire backbone with full finetune under SN-Netv2 in Table 3 of the main paper, where it achieves faster training speed while obtaining numerous complexity-performance trade-offs. In this section, we report the detailed performance curve. As shown in Figure 16 (a), with SN-Netv2, we achieve smooth performance-efficiency curve under PEFT with LoRA, which indicates that SN-Netv2 is compatible with PEFT techniques. It is also worth noting that compared to fully finetuning the entire backbone, PEFT has a noticeable performance drop at the high-FLOPs range. To this end, our experiments adopt full finetune on the entire backbone by default.
Appendix G Benchmarking Other Pretrained ViT Weights
In Figure 16 (b), we report the results of stitching different pretrained weights based on the base and large variants of ViTs, including MAE [21], SAM [26], AugReg [41] and BEiTv2 [34]. Overall, SN-Netv2 consistently generates good Pareto frontiers for these weights. As pre-training objectives and datasets differ across ViTs, our results in Figure 16 (b) demonstrate varying performance when stitching different ViT weights, where DeiT3 achieves the best performance. Therefore, we choose DeiT3 as the default weights. The details of the benchmarked ViT weights are listed as following.
Appendix H Object Detection and Instance Segmentation
Implementation details.
We experiment SN-Netv2 on COCO-2017 and adopt Mask R-CNN based ViTDet [28]. We train all models including individual anchors on 8 A100 GPUs with a total batch size of 8 for 100 epochs. We set the rank as 16 for LoRA SL. Besides, we adopt the same layer decay rate as the baselines for different anchors. All other hyperparameters adopt the default setting in detectron2 [47].
Results.
Based on DeiT3-S and DeiT3-L, we report the performance of SN-Netv2 on object detection and instance segmentation in Figure 17. Overall, we found SN-Netv2 exhibits strong flexibility on the detection task as well, as evidenced by the smooth metrics under various resource constraints. This again demonstrates that SN-Netv2 can serve as a flexible backbone on a wide range of CV tasks. In particular, under the similar training cost ($\sim$1500 GPU hours), SN-Netv2 can achieve comparable performance at the FLOPs of individually trained DeiT3-L (51.2 vs. 51.9 bbox AP), while supporting many FLOPs-accuracy trade-offs at runtime without additional training cost. However, we observe a performance gap between individually trained DeiT3-S with that of SN-Netv2 at the same FLOPs (42.9 vs. 46.6 bbox AP). For this, we hypothesize that with the heavy decoder in ViTDet (27×27$\times$27 × larger than SETR decoder), simultaneously ensuring the performance of hundreds of stitches as backbones can be more difficult. Nevertheless, SN-Netv2 still achieves superior advantage in terms of the training efficiency as it only requires training once while covering many FLOPs-accuracy trade-offs. We leave the improvement for future work.
Appendix I Memory Cost Analysis
Intuitively, SN-Netv2 will slightly increase the memory consumption as it loads two models in GPU simultaneously. However, in practice, the additional memory overhead by loading a small ViT can be relatively minor since the main bottleneck still lies in the large ViT. For example, based on COCO-Stuff-10K and a batch size of 2, SN-Netv2 by stitching DeiT3-S and DeiT3-L results in 2,432MB peak GPU memory consumption during training, which is only 1.1$\times$ higher than individual DeiT3-L training (2,226MB). In SN-Netv2, this minor memory overhead brings a favorable trade-off for obtaining hundreds of architectures.